OpenStudy (anonymous):

After a study was completed at a local college, it was found that 386 in 1,112 students live with their parents. What percent of the students live with their parents? Round to the nearest hundredth of a percent where necessary. A. 3.44% B. 34.71% C. 49.36% D. 65.28%

OpenStudy (anonymous):

Read the question carefully. What percent of students live with their parents? Well... 386 out of 1,112 students live with their parents! 386/1112. 34.71%

OpenStudy (mathstudent55):

To find what percentage a part is of the total, do this: \(\dfrac{part}{total} \times 100\% \) In your case it's: \( \dfrac{386}{1112} \times 100\% = 0.347122... \times 100\% \approx 34.71\%\)